THEORETICAL STUDY OF Er^{3+} DOPED IN SEMICONDUCTOR GaN BASED ON RELATIVISTIC QUAN-TUM CHEMICAL METHOD

\underline{YANG YANG} and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The 1.54 μm photoluminescence (^{4}I_{13/2} →^{4}I_{15/2} intra-4f shell emissions) of Er^{3+} doped in the semiconductor GaN has attracted interest for a long time, because of the potential application in optical communications due to its efficient transmission and temperature stability in semiconductor-based fibers. In our current research, we use ab initio spin-orbit configuration interaction calculations with relativistic effective core potentials to study the wave function character and energy of both the ground and excited states of ErN_{4}^{9−} aiming to compare the calculated transition frequency between those states with the available experimental data. In this calculation, ErN_{4}^{9−} has a local tetrahedral symmetry in the GaN crystal. The transition moments will be computed as well to understand the interactions giving a non-zero value.